OpenStudy (anonymous):

radical (xy)= 9 + x^2y find y' by implicit differentiation

OpenStudy (amistre64):

wht do you get on your attempt?

OpenStudy (amistre64):

i see chain rules and product rules for the most part ...

OpenStudy (freckles):

I think he can do the calculus part. I think the algebra is what scares him.

OpenStudy (anonymous):

1/(2 radical x -4x radical x)

OpenStudy (amistre64):

sqrt(xy)= 9 + x^2y (x'y+xy')/(2sqrt(xy))= 2xx' y + x^2 y' but x' wrt.x is dx/dx = 1 (y+xy')/(2sqrt(xy))= 2x y + x^2 y' looking for more of the inner workings

OpenStudy (amistre64):

to clean up . let 2 sqrt(xy) = u y+xy'= 2ux y + ux^2 y' xy'- ux^2 y'= 2ux y - y y' (x- ux^2)= y(2ux - 1) etc ...

OpenStudy (anonymous):

d/dx radical(xy) = d/dx 9 + d/dx x^2 y 1/2 radical (xy) . y' = 0 + x^2 y' + 2xy

OpenStudy (amistre64):

youve missed the chain rule on the left side the derivative of (xy) pops out

OpenStudy (amistre64):

recall:\[\sqrt{u}=u^{1/2}=1/2 u^{-1/2}~u'\] but u =xy therefore u' = x'y+xy' by the product rule

OpenStudy (anonymous):

I don't get this part

OpenStudy (anonymous):

derivative of radical(xy)

OpenStudy (aum):

\[ \frac{d}{dx}\sqrt{xy} = \frac{d}{dx}(xy)^{1/2} = \frac 12 * (xy)^{-1/2} * \frac{d}{dx}(xy) = \\ \frac 12 * (xy)^{-1/2} * (xy' + y) \]

OpenStudy (amistre64):

you are aware of the chain rule, i can see that. so to simplify it, let the innards of the radical be substituted by some function, u what is the derivative of sqrt(u) wrt.x ?

OpenStudy (anonymous):

oh i have to use the product rule to find the derivative of xy

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

is it ( 4xy radical x - 4 ) / 2x^3 radical x